# What Linguistic Features and Languages are Important in LLM Translation?

### 21 Feb 2024 v1

@@ 3.2. Language Proximity Analysis
It is evident that when including every inllama languages, syntactic proximity with English often ranks among the top features that affects translation quality. This is also evident when considering languages from other genus; English demonstrates the most substantial syntactic correlation, although there are other languages, such as Swedish and Vietnamese, that also display some degree of correlation. However, English is often not in first place when considering languages by genus (i.e., Germanic, Slavic, and Romance) despite having the highest amount of training data. Similar to when we observe that syntactic proximities to Norwegian and Catalan have higher correlations with MT scores than syntactic proximity to English when considering only outllama languages, this phenomenon is accentuated when calculating correlations by language genus. Among Germanic languages themselves, syntactic proximity to English surprisingly shows little to no correlations with either BLEU or COMET-22 scores. Instead, Germanic languages’ MT scores appear to correlate more with syntactic proximities to Dutch, Swedish, Catalan, and Bulgarian. This is also observed with Slavic languages where the MT scores generally correlate with syntactic proximities to Germanic and Romance languages except English. The syntactic proximity to English has the lowest correlation of -0.22 with BLEU and no correlation with COMET-22 scores of the Slavic languages.
Finally, when focusing exclusively on Romance languages, it is interesting to observe that proximities to languages situated on the right side of the heatmaps i.e. other genus, exhibit higher correlations with the MT scores of Romance languages, where they show no correlation with other language subsets.
@@ 4. Related Work
Our work is in line with previous studies that assess LLMs for machine translation, resembling the work conducted by Hendy et al. (2023) and Robinson et al. (2023). We extend those evaluations by investigating the influence of the languages included in the model. This aspect of the analysis was previously unattainable in the two aforementioned papers, mainly due to the lack of transparency of the data used for training LLMs. Our method of analysis is closely related to the one included in the work of Robinson et al. (2023) where languages of interest are represented as feature vectors to do further evaluation. Our objective is to extend that exploration by encompassing other linguistic features obtained from the URIEL typological database Littell et al. (2017). We are interested in the phenomenon observed in the work of Lin et al. (2019) which also uses distances retrieved from URIEL Typological Database Littell et al. (2017). This work shows that however important dataset statistics features are compared to linguistic features, there are still cases where using them alone results in poor performance. This phenomenon drives us to conduct a more comprehensive exploration of linguistic features.
@@ 5. Conclusion
We provide a comprehensive evaluation of machine translation task in Llama2 in its seen and unseen languages. In this work, we provide machine translation scores for 26 non-English languages covered in the Llama2 model training data using the Llama2-7B model. We also added 15 additional languages not in the Llama2 training data using the 7B, 7B-chat, 13B, and 13B-chat versions. Our results show that Llama2 is capable of translating into languages it is unfamiliar with, although this phenomenon is observed only for some languages. We also demonstrate that the model scale has the most substantial impact on translation performance, while the impact of using the chat versions of Llama2 and adding the shot count varies by language. We calculate and visualize the correlations between language feature distances and machine translation scores. Our analysis reveals that syntactic similarity is not always the predominant linguistic factor influencing machine translation scores. Furthermore, despite English having the most training data, there are often other languages, whose features consistently exhibit equally or even stronger correlations in determining machine translation scores albeit having much fewer training data. Our findings pose a unique perspective on state-of-the-art LLMs, suggesting that the prevailing focus on English-centric models may not be the most optimal approach, potentially opening the door to more effective multilingual systems that could be shaped by languages other than English.
@@
Our research heavily depends on the language distances obtained from the URIEL typological database, as introduced by Littell et al. (2017). The original authors noted that many languages in the database may have missing features, which means the accuracy of our findings is constrained by the methods used to estimate or compensate for these missing features. Our evaluation with the COMET-22 metric is only done for languages supported in their models. However, we are aware that the model may not be equally reliable for all languages, thus the reliability of the COMET-22 heatmap goes as far as the COMET-22 model.
In an ideal scenario, it would be advantageous to include all languages from the FLORES-200 benchmark and all available versions of Llama2 to provide more evidence of the effectiveness of model scale and the overall generalizability of our findings. Unfortunately, our research is constrained by limited computational resources, preventing us from achieving this comprehensive coverage. We exclude to-English translation directions as Llama2 is likely to be trained on English Wikipedia corpora. We also exclude prompting languages in outllama using various dictionary-based prompting techniques due to the challenging work required to collect accurate dictionary entries for low-resource languages. However, we leave this for future work.
@@ 6. Bibliographical References
- Guzmán et al. (2019) Francisco Guzmán, Peng-Jen Chen, Myle Ott, Juan Pino, Guillaume Lample, Philipp Koehn, Vishrav Chaudhary, and Marc’Aurelio Ranzato. 2019. The FLORES evaluation datasets for low-resource machine translation: Nepali–English and Sinhala–English. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP), pages 6098–6111, Hong Kong, China. Association for Computational Linguistics.

### 05 Apr 2024 v2

@@ 3.2. Language Importance Analysis
When considering languages from other genera English demonstrates the most substantial syntactic correlation with MT performance, although there are other languages, such as Swedish and Vietnamese, that also display some degree of correlation. However, despite having the highest amount of training data, English is often not in the first place when considering languages by genus (e.g., Germanic, Slavic, and Romance). Similar to when we observe that syntactic proximity to Norwegian and Catalan have higher correlations with MT scores than syntactic proximity to English when considering only outllama languages, this phenomenon is accentuated when calculating correlations by genus. Among Germanic languages, syntactic proximity to English surprisingly shows little to no correlation with either BLEU or COMET-22 scores. Instead, Germanic languages’ MT scores appear to correlate more with syntactic proximity to Dutch, Swedish, Catalan, and Bulgarian. This is also observed in Slavic languages where the MT scores generally correlate with syntactic proximity to most Germanic and Romance languages except English. With Slavic languages, syntactic proximity to English has the lowest correlation on BLEU and almost no correlation on COMET-22 scores. Finally, when focusing exclusively on Romance languages, it is interesting to observe that proximities to languages situated on the right side of the heatmaps i.e. other genera, exhibit higher correlations while they show no correlation when only considering other language subsets (Figure 2 and 3).
@@ 4. Related Work
Our work aligns with previous studies that assess LLMs for translation, resembling the work by Hendy et al. (2023) and Robinson et al. (2023). We aim to extend such evaluations further by investigating the influence of the languages included in the training data of the model, which was previously underexplored due to the lack of transparency of LLMs used. Our method of analysis, similar to the work of Robinson et al. (2023), investigates feature importance. Our objective is to extend that exploration by encompassing other linguistic features obtained from the URIEL typological database Littell et al. (2017). We are interested in the phenomenon observed in the work of Lin et al. (2019) which shows that however important dataset statistics are compared to linguistic features, there are cases where using them alone to choose transfer languages results in poor performance. This phenomenon drove us to conduct a more comprehensive exploration of linguistic features.
@@ 5. Conclusion
We provide a comprehensive evaluation of machine translation in Llama2 for languages seen or unseen in its training data. In this work, we provide EnglishX machine translation scores of Llama2 7B for 26 languages reported to be in the training data of Llama2 models. We also evaluated 15 additional languages that are not reported to be in Llama2 training data using the 7B, 7B-chat, 13B, and 13B-chat Llama2 models. Our results show that Llama2 is capable of translating into languages it is unfamiliar with, although this phenomenon is observed only in some languages. We demonstrate that model scaling has the most substantial impact when compared to instruction tuning and adding shot count, whose improvements vary by language. We also modeled the linear relationship of linguistic distances and translation quality through correlation scores and revealed that syntactic similarity is not the only feature that displays strong correlations with machine translation scores. Furthermore, despite English having the most training data, there are other languages (e.g. Swedish, Catalan) whose linguistic distances exhibit comparable correlation scores to English albeit having much fewer training data. Our findings pose a unique perspective on the current landscape of language models, suggesting that the prevailing focus on English-centered models may not be the most optimal setup for multilingual models. We hope to open doors toward more effective and training-data-efficient multilingual systems that are shaped by languages other than English, thus promoting digital language equality and sustainability.
@@
Our research heavily depends on the language distances obtained from the URIEL typological database, as introduced by Littell et al. (2017). The original authors noted that many languages in the database may have missing features, which means the accuracy of our findings is constrained by the methods used to compensate for these missing features. Our evaluation with the COMET-22 metric is only done for languages supported in their models. However, the model may not be equally reliable for all languages, thus the COMET-22 correlations are only as accurate as the COMET-22 model. Furthermore, there are other ways to model the relationship between language feature distances and machine translation scores. We leave such investigations for future work. We also left out positively correlated features in our analysis as they are not readily interpretable in the context of our analysis.
In an ideal scenario, it would be advantageous to include all languages from the FLORES-200 benchmark and all available versions of Llama2 and other multilingual models to provide more evidence of the effectiveness of scaling parameter count and the overall generalizability of our findings. Unfortunately, our research is constrained by limited computational resources, preventing us from achieving this comprehensive coverage. We exclude XEnglish translation directions as Llama2 is likely trained on English Wikipedia. We also exclude prompting languages in outllama using various dictionary-based prompting techniques due to the challenging work required to collect accurate dictionary entries for low-resource languages. However, we leave this for future work.
@@
Acknowledgements
Authors from Indonesian institutions are supported by the Indonesian Ministry of Education, Culture, Research, and Technology (MoECRT) ACE Open Research program. This work is also supported in part by the Indonesia-US Research Collaboration in Open Digital Technology grant funded by the Indonesian Ministry of Education, Culture, Research, and Technology. The authors would like to thank the Indonesian government for their funding and Boston University for providing essential computing resources through the Shared Computing Cluster (SCC).
@@ 6. Bibliographical References
- Brown et al. (2020) Tom B. Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel M. Ziegler, Jeffrey Wu, Clemens Winter, Christopher Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei. 2020. Language models are few-shot learners.
@@
- Strubell et al. (2019) Emma Strubell, Ananya Ganesh, and Andrew McCallum. 2019. Energy and policy considerations for deep learning in NLP. In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, pages 3645–3650, Florence, Italy. Association for Computational Linguistics.
@@
- Wijaya et al. (2017) Derry Tanti Wijaya, Brendan Callahan, John Hewitt, Jie Gao, Xiao Ling, Marianna Apidianaki, and Chris Callison-Burch. 2017. Learning translations via matrix completion. In Proceedings of the 2017 Conference on Empirical Methods in Natural Language Processing, pages 1452–1463, Copenhagen, Denmark. Association for Computational Linguistics.